OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). A. y = -2x - 4 B. y = -2x - 8 C. y = -2x + 18 D. y = -2x – 12

OpenStudy (kenljw):

y = -2 x + b -8 = -2(2) + b b = -4 y = -2 x - 4